if the MPC is an economy is .8, government could shift the aggregate demand curve rightward by $100 billion by:
Answer #1

ANSWER:

TAX MULTIPLIER = -MPC / (1 - MPC) = -0.8 / (1 - 0.8) = -0.8 / 0.2 = -4

Tax multiplier = increase in income / decrease in tax

-4 = $100 billion / decrease in tax

decrease in tax = $100 billion / -4 = -$25 billion

so the decrease int ax by $25 billion.